Solve the system by substitution.
4x – 4y = –4
y = 2x + 0

Answer by rapaljer(4671)

4x – 4y = –4
y = 2x + 0

Since the second equation is in the form y =___, it will be convenient to substitute this back into the first equation!

4x - 4y = -4
4x-4(2x) = -4
4x - 8x = -4
-4x = -4
x=1

Now, substitute this into either equation. The easiest was is to use the y = __ equation.

y=2x
y=2(1)
y=2

The solution is (1,2).

Check in the OTHER equation:
4x-4y =-4
4(1)-4(2)=-4
4-8=-4 It checks!!

R^2 at SCC